# A Heuristic Algorithm Based on Beam Search and Iterated Local Search for the Maritime Inventory Routing Problem

### v1

@@ 6 Conclusions
The main objective of this work was to develop a heuristic approach capable of solving a MIRP without relying on mathematical optimization models, and therefore (commercial) MILP solvers, while also aiming to obtain high-quality solutions for MIRPLib Group 2 instances. To this end, we put forward a method that combines BS with ILS. The proposed BS strategy enabled the exploration of a restricted subset of promising solution paths, significantly reducing computational effort while directing the search toward high-quality solutions. This was made possible due the implementation of a newly proposed greedy algorithm, which played a crucial role in efficiently completing partial solutions during BS. Furthermore, (iterated) local search procedures demonstrated to be essential to refine and improve the quality of solutions, producing competitive results. In particular, comparative analysis with MIRPLib benchmarks revealed ten new best-known solutions.
The proposed methodology is flexible enough to easily adjust its performance by modifying the number of nodes in the BS. This allows for a controlled trade-off between smaller gaps and higher computational times, or vice versa. For smaller instances, the methodology shows strong efficiency, reaching solutions with low optimality gaps within minutes, and in some cases surpassing the quality of known benchmark results. For larger instances, maintaining a low gap requires increasing computational time (through the number of nodes in the BS). Nevertheless, satisfactory results can still be achieved in less than 24 hours.
Future work includes the development of new local search operators and acceleration mechanisms to speed up the convergence of the method. Furthermore, adaptive perturbation mechanisms could be designed to improve the diversification phase of the proposed algorithm, helping it escape local optima.
7 Comparison of Results: BS versus Greedy Heuristic
@@
Instance Bound ObjVal After BS After LS After ILS GAP ObjVal (%)
LR1_DR02_VC01_V6a 180 52166.00 52167.00 52642.82 52167.21 52167.21 0.00
360 107994.00 108141.00 129438.62 108141.00 108141.00 0.00
LR1_DR02_VC02_V6a 180 128106.00 129372.00 136876.18 130603.20 130603.20 0.95
360 275347.00 283031.00 327394.47 300509.56 298807.87 5.57
LR1_DR02_VC03_V7a 120 39318.00 40446.00 49126.32 43045.13 43018.93 6.36
LR1_DR02_VC03_V8a 120 43717.00 43721.00 59676.14 47892.73 47508.72 8.66
180 66989.00 68153.00 94931.65 75464.21 73314.49 7.57
LR1_DR02_VC04_V8a 120 41277.00 41657.00 47211.03 43183.50 42206.12 1.32
LR1_DR02_VC05_V8a 120 36088.00 36659.00 38024.22 37255.92 37008.85 0.95
LR1_DR03_VC03_V13b 120 118706.00 124921.00 151117.90 129835.53 128395.38 2.78
LR1_DR03_VC03_V16a 120 71635.00 82837.00 171093.48 153981.07 116603.27 40.76
LR1_DR04_VC03_V15a 120 72108.00 73312.00 96851.16 76859.70 74896.02 2.16
LR1_DR04_VC03_V15b 120 102148.00 117812.00 153494.96 130322.79 123931.88 5.19
LR1_DR04_VC05_V17a 120 71572.00 72876.00 80193.56 74921.30 74119.90 1.71
LR1_DR04_VC05_V17b 120 84970.00 105766.00 146881.76 123865.09 116463.71 10.11
@@
Instance - Horizon Cost Max Cost Mean Cost Median
LR1_DR02_VC01_V6a_120 33923.34 33939.33 34266.97
LR1_DR02_VC01_V6a_180 52647.27 52481.66 52601.11
LR1_DR02_VC01_V6a_360 109414.31 110131.37 109655.42
LR1_DR02_VC02_V6a_120 77395.06 76818.15 76624.52
LR1_DR02_VC02_V6a_180 135993.47 135549.76 135545.76
LR1_DR02_VC02_V6a_360 304937.43 303904.98 302975.50
LR1_DR02_VC03_V7a_120 42910.78 42464.15 42991.15
LR1_DR02_VC03_V7a_180 67109.18 67550.13 67528.97
LR1_DR02_VC03_V7a_360 140255.62 143311.20 142994.77
LR1_DR02_VC03_V8a_120 47900.23 48310.53 48050.30
LR1_DR02_VC03_V8a_180 77256.03 78405.74 77311.96
LR1_DR02_VC03_V8a_360 174802.30 175016.07 173030.24
LR1_DR02_VC04_V8a_120 42091.77 42164.18 42218.00
LR1_DR02_VC04_V8a_180 68388.41 67871.17 67815.13
LR1_DR02_VC04_V8a_360 145597.41 145606.29 145889.97
LR1_DR02_VC05_V8a_120 37002.98 36954.36 36988.57
LR1_DR02_VC05_V8a_180 58923.26 58750.84 58762.10
LR1_DR02_VC05_V8a_360 124037.47 123887.84 123824.35
LR1_DR03_VC03_V10b_120 110858.34 111754.60 114018.19
LR1_DR03_VC03_V10b_180 145883.67 145080.40 146118.27
LR1_DR03_VC03_V10b_360 370964.74 362337.30 358205.90
LR1_DR03_VC03_V13b_120 136033.99 134909.53 135552.79
LR1_DR03_VC03_V13b_180 173509.93 172567.81 172606.80
LR1_DR03_VC03_V13b_360 341677.79 341456.29 341661.21
LR1_DR03_VC03_V16a_120 138716.00 137684.78 131597.83
LR1_DR03_VC03_V16a_180 276705.77 276930.69 274921.64
LR1_DR03_VC03_V16a_360 682519.75 682040.67 682703.66
LR1_DR04_VC03_V15a_120 82961.41 85314.85 84632.05
LR1_DR04_VC03_V15a_180 148329.62 148596.19 148713.50
LR1_DR04_VC03_V15a_360 301765.71 301804.37 301325.94
LR1_DR04_VC03_V15b_120 143300.31 143387.59 143254.13
LR1_DR04_VC03_V15b_180 212916.42 212891.02 212375.42
LR1_DR04_VC03_V15b_360 387058.14 382095.79 381325.83
LR1_DR04_VC05_V17a_120 75673.42 75783.27 75735.27
LR1_DR04_VC05_V17a_180 122075.49 121668.81 122024.03
LR1_DR04_VC05_V17a_360 261484.75 260646.48 260256.08
LR1_DR04_VC05_V17b_120 134376.88 133850.50 132773.17
LR1_DR04_VC05_V17b_180 197886.09 198980.83 199055.59
LR1_DR04_VC05_V17b_360 412700.92 408949.66 409822.29
8 Complete Final Results
This appendix provides the complete final results from the computational experiments conducted with the proposed heuristic. Detailed results of total cost and time are shown for each instance, in three different time horizons for different values of $N$.
Table 11: 120 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)
@@ 8 Complete Final Results
10 400.556,07 5.433,52 315.584,26 1.427,82 304.041,16 49.687,21
100 354.136,74 63.738,30 304.932,19 13.348,48
Table 12: 180 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)
@@
10 667.435,32 10.885,22 511.421,68 4.169,28
Table 13: 360 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)

### v2

@@ 6 Conclusions
The main objective of this work was to develop a heuristic approach capable of solving a MIRP without relying on mathematical optimization models and therefore (commercial) MILP solvers, while also aiming to obtain high-quality solutions for MIRPLib Group 2 instances. To this end, we propose a method that combines BS with ILS. The proposed BS strategy allowed the exploration of a restricted subset of promising solution paths, significantly reducing computational effort while directing the search toward high-quality solutions. This was possible due to the implementation of a greedy algorithm, which played a crucial role in efficiently completing partial solutions during BS. Furthermore, (iterated) local search procedures demonstrated to be essential to refine and improve the quality of solutions, producing competitive results. In particular, a comparative analysis with MIRPLib benchmarks presented 19 new best-known solutions.
The proposed methodology is flexible enough to adjust its performance by modifying the number of nodes in the BS. This allows for a controlled trade-off between smaller gaps and higher computational times, or vice versa. For smaller instances, the methodology shows strong efficiency, reaching solutions with low optimality gaps within minutes, and in some cases surpassing the quality of known benchmark results. For larger instances, maintaining a low gap requires increasing computational time (through the number of nodes in the BS). Nevertheless, satisfactory results can still be achieved in less than 24 hours.
Future work includes the development of new local search operators and acceleration mechanisms to speed up the convergence of the method. In addition, adaptive perturbation mechanisms could be designed to improve the diversification phase of the proposed algorithm, helping it escape local optima.
Acknowledgments
This work was partially supported by CAPES (Coordenacão de Aperfeiçoamento de Pessoal de Níıvel Superior - Finance Code 001), CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico - projects 315361/2020-4, 422470/2021-0, 406245/2021-5, 309580/2021-8 and 314420/2023-1), FAPESQ (Paraíba State Research Foundation - project 041/2023) and by FAPERJ (Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro - projects E-26/201.417/2022, E-26/010.002232/2019 and E-26/210.041/2023).
Appendix A Comparison of Results: BS versus Greedy Heuristic
@@
Instance Horizon Bound Obj After BS After LS After ILS Gap Obj
LR1_DR02_VC01_V6a 180 52,166.00 52,167.00 52,642.82 52,167.21 52,167.21 0.00%
360 107,994.00 108,141.00 129,438.62 108,141.00 108,141.00 0.00%
LR1_DR02_VC02_V6a 180 128,106.00 129,372.00 136,876.18 130,603.20 130,603.20 0.95%
360 275,347.00 283,031.00 327,394.47 300,509.56 298,807.87 5.57%
LR1_DR02_VC03_V7a 120 39,318.00 40,446.00 49,126.32 43,045.13 43,018.93 6.36%
LR1_DR02_VC03_V8a 120 43,717.00 43,721.00 59,676.14 47,892.73 47,508.72 8.66%
180 66,989.00 68,153.00 94,931.65 75,464.21 73,314.49 7.57%
LR1_DR02_VC04_V8a 120 41,277.00 41,657.00 47,211.03 43,183.50 42,206.12 1.32%
LR1_DR02_VC05_V8a 120 36,088.00 36,659.00 38,024.22 37,255.92 37,008.85 0.95%
LR1_DR03_VC03_V13b 120 118,706.00 124,921.00 151,117.90 129,835.53 128,395.38 2.78%
LR1_DR03_VC03_V16a 120 71,635.00 82,837.00 171,093.48 153,981.07 116,603.27 40.76%
LR1_DR04_VC03_V15a 120 72,108.00 73,312.00 96,851.16 76,859.70 74,896.02 2.16%
LR1_DR04_VC03_V15b 120 102,148.00 117,812.00 153,494.96 130,322.79 123,931.88 5.19%
LR1_DR04_VC05_V17a 120 71,572.00 72,876.00 80,193.56 74,921.30 74,119.90 1.71%
LR1_DR04_VC05_V17b 120 84,970.00 105,766.00 146,881.76 123,865.09 116,463.71 10.11%
@@
Instance Horizon Minimum Cost Mean Cost Median Cost
LR1_DR02_VC01_V6a 120 33,923.34 33,939.33 34,266.97
LR1_DR02_VC01_V6a 180 52,647.27 52,481.66 52,601.11
LR1_DR02_VC01_V6a 360 109,414.31 110,131.37 109,655.42
LR1_DR02_VC02_V6a 120 77,395.06 76,818.15 76,624.52
LR1_DR02_VC02_V6a 180 135,993.47 135,549.76 135,545.76
LR1_DR02_VC02_V6a 360 304,937.43 303,904.98 302,975.50
LR1_DR02_VC03_V7a 120 42,910.78 42,464.15 42,991.15
LR1_DR02_VC03_V7a 180 67,109.18 67,550.13 67,528.97
LR1_DR02_VC03_V7a 360 140,255.62 143,311.20 142,994.77
LR1_DR02_VC03_V8a 120 47,900.23 48,310.53 48,050.30
LR1_DR02_VC03_V8a 180 77,256.03 78,405.74 77,311.96
LR1_DR02_VC03_V8a 360 174,802.30 175,016.07 173,030.24
LR1_DR02_VC04_V8a 120 42,091.77 42,164.18 42,218.00
LR1_DR02_VC04_V8a 180 68,388.41 67,871.17 67,815.13
LR1_DR02_VC04_V8a 360 145,597.41 145,606.29 145,889.97
LR1_DR02_VC05_V8a 120 37,002.98 36,954.36 36,988.57
LR1_DR02_VC05_V8a 180 58,923.26 58,750.84 58,762.10
LR1_DR02_VC05_V8a 360 124,037.47 123,887.84 123,824.35
LR1_DR03_VC03_V10b 120 110,858.34 111,754.60 114,018.19
LR1_DR03_VC03_V10b 180 145,883.67 145,080.40 146,118.27
LR1_DR03_VC03_V10b 360 37,0964.74 362,337.30 358,205.90
LR1_DR03_VC03_V13b 120 136,033.99 134,909.53 135,552.79
LR1_DR03_VC03_V13b 180 173,509.93 172,567.81 172,606.80
LR1_DR03_VC03_V13b 360 341,677.79 341,456.29 341,661.21
LR1_DR03_VC03_V16a 120 138,716.00 137,684.78 131,597.83
LR1_DR03_VC03_V16a 180 276,705.77 276,930.69 274,921.64
LR1_DR03_VC03_V16a 360 682,519.75 682,040.67 682,703.66
LR1_DR04_VC03_V15a 120 82,961.41 85,314.85 84,632.05
LR1_DR04_VC03_V15a 180 148,329.62 148,596.19 148,713.50
LR1_DR04_VC03_V15a 360 301,765.71 301,804.37 301,325.94
LR1_DR04_VC03_V15b 120 143,300.31 143,387.59 143,254.13
LR1_DR04_VC03_V15b 180 212,916.42 212,891.02 212,375.42
LR1_DR04_VC03_V15b 360 387,058.14 382,095.79 381,325.83
LR1_DR04_VC05_V17a 120 75,673.42 75,783.27 75,735.27
LR1_DR04_VC05_V17a 180 122,075.49 121,668.81 122,024.03
LR1_DR04_VC05_V17a 360 261,484.75 260,646.48 260,256.08
LR1_DR04_VC05_V17b 120 134,376.88 133,850.50 132,773.17
LR1_DR04_VC05_V17b 180 197,886.09 198,980.83 199,055.59
LR1_DR04_VC05_V17b 360 412,700.92 408,949.66 409,822.29
Average 169,179.11 168,765,36 168,403.956
Appendix B Complete Final Results
This appendix provides the complete results from the computational experiments conducted with the proposed heuristic. Detailed results of the average total cost and time are shown for each instance, in three different time horizons for different values of $N$.
After BS After LS After ILS
Instance Cost Time(s) Cost Time(s) Cost Time(s)
@@ Appendix B Complete Final Results
10 400,556.07 5,433.52 315,584.26 1,427.82 304,041.16 49,687.21
100 354,136.74 63,738.30 304,932.19 13,348.48
Table 11: Results for instance with horizon = 120
After BS After LS After ILS
Instance Cost Time(s) Cost Time(s) Cost Time(s)
@@
10 667,435.32 10,885.22 511,421.68 4,169.28
Table 12: Results for instance with horizon = 180
After BS After LS After ILS
Instance Cost Time(s) Cost Time(s) Cost Time(s)